OpenStudy (anonymous):

Solve the following system of equations. x + y = 3 2x + 2y = 6 (Points : 4) (0, 3) (1.5, 1.5) no solution infinite number of solutions could someone explain how i find the answer?

OpenStudy (dtan5457):

how would you like to solve this? substitution, graphically <easiest, elimination? choose.

OpenStudy (dtan5457):

@marie123

OpenStudy (anonymous):

i dont know???

Nnesha (nnesha):

2 answer choices r order pair so you can plug x and y (x , y) values plug back into the equation if you get both side equal then that is your right answer

OpenStudy (dtan5457):

ok well say you choose substitution. you have the equations x+y=3 2x+2y=6 get one of the variables alone like for example y y=3-x plug that into second equation 2x+2(3-x)=6 2x+6-2x=6 -6 -6 x=0 now you can easily get y by solving any of the equations knowing x=0 x+y=3 0+y=3 y=3